OpenStudy (anonymous):

Given that theta=tan^-1(4/3), find the exact values of sin(theta), cos(theta), cot(theta), sec(theta), and csc(theta).

myininaya (myininaya):

\[\theta=\tan^{-1}(4/3)=>\tan(\theta)=\frac{4}{3}=\frac{opp}{adj}\]